OpenStudy (anonymous):

Hi..I have a final on Wednesday and I have a question about system of equation: x-y=-1 x^2-y=5 Please help! Thanks!

OpenStudy (anonymous):

Whats the question? Solve using substitution and factor. You'll have a lot of answers cause of +,- when you sqrt.

OpenStudy (anonymous):

take value of x=-1+y and put in second equation

OpenStudy (anonymous):

Thank you guys! I will try again! :)

OpenStudy (anonymous):

\[x-y=-1\iff x+1=y\]\[ x^2-y=5 \] \[x^2-(x+1)=5\]might be easiest

OpenStudy (anonymous):

Alternate method is to subtract the two equations. That will allow the y values to cancel, and leave you with a quadratic in x. Find the roots, then substitute to find the y values.

OpenStudy (anonymous):

The answers at the back of the review say: (-2,-1), (3,4)

OpenStudy (anonymous):

I like that! Thank you AnimalAin :)

OpenStudy (anonymous):

No sweat. Make sure you are careful with the signs, and it will come out to the same quadratic Sattellite73 has directly above my post.

OpenStudy (anonymous):

Thank you everybody! you are the best!

OpenStudy (anonymous):

I think I got it..

OpenStudy (precal):

Many ways to solve systems, pick a few that you like and study those. I like to use elimination method first, graphing method second, substitution when I have to.

OpenStudy (anonymous):

aww..Thank you good advice! :)

OpenStudy (precal):

Always check your solution, no matter what method you use, it has to work for both equations or (all)

OpenStudy (precal):